# A Survey on Efficient Vision-Language-Action Models

### v1

@@ Abstract
Vision-Language-Action models (VLAs) represent a significant frontier in embodied intelligence, aiming to bridge digital knowledge with physical-world interaction. While these models have demonstrated remarkable generalist capabilities, their deployment is severely hampered by the substantial computational and data requirements inherent to their underlying large-scale foundation models. Motivated by the urgent need to address these challenges, this survey presents the first comprehensive review of Efficient Vision-Language-Action models (Efficient VLAs) across the entire data-model-training process. Specifically, we introduce a unified taxonomy to systematically organize the disparate efforts in this domain, categorizing current techniques into three core pillars: (1) Efficient Model Design, focusing on efficient architectures and model compression; (2) Efficient Training, which reduces computational burdens during model learning; and (3) Efficient Data Collection, which addresses the bottlenecks in acquiring and utilizing robotic data. Through a critical review of state-of-the-art methods within this framework, this survey not only establishes a foundational reference for the community but also summarizes representative applications, delineates key challenges, and charts a roadmap for future research. We maintain a continuously updated project page to track our latest developments: https://evla-survey.github.io/
@@ 1 Introduction
Vision-Language-Action Models (VLAs) herald a breakthrough in embodied intelligence, forging a direct pathway from digital cognition to physical-world actuation. Building upon large-scale pretrained vision-language models (VLMs), representative VLAs such as OpenVLA [1], $\pi_{0}$ [2], and others [3, 4] achieve multimodal perception and understanding by fusing visual and linguistic modalities, thereby enabling effective action generation for real-world interactions. Their applications are expansive and transformative, spanning autonomous driving [5, 6], industrial manufacturing [7], medical robotics [8], laboratory automation [9], and beyond, with the potential to revolutionize industries, enhance human-robot collaboration, and tackle foundational challenges in robotics.
Figure 1: Necessity of Efficient VLAs. This figure highlights the disparity between powerful but resource-intensive foundation VLAs and the practical deployment requirements of diverse edge robotic platforms. Bridging this gap by developing more compact, economical, and applicable solutions is the primary motivation for pursuing efficient VLAs.
At the core of VLAs lies a modular architecture that fuses multimodal inputs through a vision encoder and a language encoder, which process raw cross-modal data into high-dimensional embeddings; a large language model backbone, serving as the reasoning hub to contextualize instructions and perceptions; and an action decoder, which maps fused representations to low-level control signals for embodied execution. Current research on VLAs centers on unifying perception, language understanding, and action generation in end-to-end frameworks, with a strong emphasis on architectural innovations, large-scale multimodal pretraining, and strategies to enhance generalization across diverse tasks, embodiments, and environments.
@@
Figure 2: The Organization of Our Survey. We systematically categorize efficient VLAs into three core pillars: (1) Efficient Model Design , encompassing efficient architectures and model compression techniques; (2) Efficient Training , covering efficient pre-training and post-training strategies; and (3) Efficient Data Collection, including efficient data collection and augmentation methods. The framework also reviews VLA foundations, key applications, challenges, and future directions, establishing the groundwork for advancing scalable embodied intelligence.
Despite propelling embodied intelligence to unprecedented heights by seamlessly integrating perception, reasoning, and control, their efficacy remains critically hindered by profound efficiency bottlenecks. These scaling-dependent paradigms inherit the resource-intensive nature of their LLM and VLM foundations, manifesting in: (1) high inference latency and insufficient frequency, incompatible with the sub-second control cycles required for responsive manipulation; (2) prohibitive pretraining demands, with models like $\pi_{0}$ [2] requiring 10,000 hours of robotic trajectories and OpenVLA [1] consuming 21,500 A100-GPU hours across a 64-GPU cluster; and (3) time-consuming and labor-intensive large datasets, rendering training infeasible in resource-constrained environments. As illustrated in  Fig.˜1, foundational VLA models grapple with significant efficiency challenges, typically characterized by the need for computing power support from cloud servers, and cannot be efficiently developed and deployed on edge devices. Addressing these through efficient VLA designs [10, 11, 12, 13, 14, 15] promises broader accessibility for resource-constrained researchers and enables deployment in diverse scenarios, while enhancing inference speed to yield smoother control and improved task performance.
@@
In this survey, we are dedicated to efficient VLAs in the embodied AI domain, aiming to systematize the rapidly evolving field by presenting a detailed taxonomy. We define efficient VLAs as systems aiming for high performance while maintaining low overhead across the development and deployment lifecycle. As illustrated in Fig.˜2, we organize existing approaches into a structured taxonomy comprising three core technical directions: (1) Efficient Model Design encompasses innovative strategies to optimize the architectural and inference efficiency of VLAs. This includes Efficient Architectures, which offers innovative structural optimizations alongside scalable frameworks, and Model Compression, which involves techniques to compress model parameters and refine token handling. (2) Efficient Training incorporates advanced approaches to lessen the computational and data burdens during the VLA training process. This includes Efficient Pre-training, which utilizes optimized data techniques and refined action representations, and Efficient Post-training that applies customized supervised fine-tuning and reinforced learning strategies. (3) Efficient Data Collection integrates cutting-edge methods to efficiently enhance data gathering and enrichment for VLA development, which employs interactive, simulated, reutilize-oriented, and self-driven techniques to acquire datasets in a more scalable and efficient way.
@@
This pioneering taxonomy marks the seminal comprehensive effort to synthesize cutting-edge advancements and technical trajectories in efficient VLAs, and is the first review of efficient VLAs that covers VLA development and deployment around the full “data-model-training” process, setting a foundation that galvanizes future research and steers the field toward a cohesive frontier. Although several excellent surveys [16, 17, 18, 19, 20, 21] have charted the landscape of general Vision-Language-Action Models or the broader field of embodied AI, a systematic review dedicated to the crucial aspect of efficiency in VLAs has been conspicuously absent. This work fills this critical void, uniquely consolidating the disparate research efforts focused on optimizing VLAs for real-world deployment in resource-constrained scenarios, making it an indispensable reference for researchers and practitioners. In doing so, this survey not only complements existing literature but also formally defines and structures the emerging subfield of efficient VLAs.
@@
  • Pioneering Survey: To the best of our knowledge, this work presents the first comprehensive survey specifically dedicated to the realm of efficient VLAs that covers the entire “data-model-training” process. It fills a critical gap in the literature and aims to establish a foundational reference for the research community.
@@
The remaining sections of this survey are organized as illustrated in Fig.˜2. We begin in section˜2 by introducing the foundational concepts of Vision-Language-Action models, delineating the computational challenges that necessitate the pursuit of efficiency. Section 3 delves into efficient model design, systematically categorizing techniques that span efficient architectural innovations and model compression. Section 4 investigates efficient training paradigms, encompassing strategies to reduce the computational overhead during both pre-training and post-training stages. Section 5 examines the crucial role of efficient data collection, exploring methodologies for scalable data collection and effective augmentation to maximize data utility. We then survey key applications in section˜6 where efficient VLAs have demonstrated significant potential, highlighting their practical utility in real-world scenarios. Section 7 discusses current challenges and outlines promising future research directions. Finally, section˜8 summarizes the survey.
@@ 2 Vision-Language-Action Models
The emergence of Vision-Language-Action (VLA) models represents a paradigm-shifting frontier in embodied AI, redefining the synergy between human-like perception and autonomous action. Classic VLA frameworks extend the prowess of pre-trained Vision-Language Models (VLMs) into actionable robotic control, seamlessly integrating multimodal perception and language to drive precise motor responses. However, their reliance on extensive computational clusters and vast, diverse training datasets often hinders development and deployment in resource-constrained settings, underscoring the urgent need for efficient VLA solutions. In response, researchers committed to Efficient VLAs are actively pursuing innovative architectures and optimization techniques to enhance efficiency in real-world applications.
@@ 2.1 An Overview of VLAs
VLAs mark a critical milestone in embodied AI, extending the capabilities of pre-trained VLMs to enable precise robotic action execution. At their core, these models tackle the fundamental task of translating multimodal sensory and linguistic inputs into coherent physical responses, addressing the gap between abstract understanding and practical motor control in dynamic environments.
@@ 2.1.1 Foundational Pipeline
LLM Backbone. After a projector $P(\cdot)$ that ensures seamless alignment and mitigates modality gaps, the multimodal sequence then feeds into the LLM backbone $LLM(\cdot)$—the center for semantic reasoning. Spanning LLaVA [26], Qwen [27], PaLM [28], Gemma [29], Llama [30], Mamba [31], VILA-U [32], and Vicuna [33], pre-trained LLMs process the fused embeddings for task planning, which can be formulated as follows:
@@
2.1.2 Datasets and Benchmarks
The advancement of VLAs is intrinsically linked to the availability of large-scale, diverse, and well-structured training data and evaluation benchmarks. The field has seen the curation of substantial datasets, which can be broadly categorized into real-world and simulation-sourced collections.
@@ 2.1.2 Datasets and Benchmarks
Real-world Datasets. A cornerstone of VLA training is the Open X-Embodiment (OXE [38]) dataset. This collaborative effort aggregates data from multiple robotics labs, encompassing a vast array of tasks, robot morphologies, and environments. Its scale and diversity are instrumental for training models with strong generalization and transfer learning capabilities. Complementing OXE are other significant datasets such as BridgeData and BridgeData V2 [39, 40], which focuses on cross-domain tasks, and DROID [41], known for its large-scale dexterous teleoperated data “in the wild”. These datasets provide the crucial “embodied” experience necessary for VLAs to learn the dynamics and constraints of the physical world.
Simulation Datasets. Due to the cost and scalability limitations of physical data collection, simulation platforms play an indispensable role in generating the large-scale training data required for VLAs. RLBench [42] offers a vast library of predefined manipulation tasks with standardized evaluation protocols, serving as a popular testbed for algorithm development and providing substantial pre-training data. Further expanding the scope of simulation, RoboCasa [43] provides a large-scale simulation environment focusing on everyday tasks for generalist robots, significantly enriching the diversity and practicality of training scenarios. Meanwhile, RoboGen [44] introduces a novel paradigm by leveraging generative simulation to automatically produce a vast array of tasks, scenes, and training demonstrations, thereby pushing towards unleashing infinite data for automated robot learning. These simulated environments collectively enable rapid prototyping, extensive data augmentation, and the safe testing of policies under controlled conditions prior to real-world deployment.
Benchmarks. Rigorous evaluation is paramount for tracking progress. A suite of benchmarks has been established to assess different aspects of VLA performance. Meta-World [45] provides a set of challenging robotic manipulation tasks, while LIBERO [46] focuses on lifelong learning and knowledge retention in sequential task settings. CALVIN [47] and SIMPLER [48] are designed to benchmark generalization to novel instructions, objects, and scenes, and long-horizon reasoning. More recently, VLABench [49] has emerged as a benchmark specifically tailored for evaluating the visual grounding and action reasoning capabilities of VLAs in embodied scenarios. The collective use of these benchmarks ensures a comprehensive assessment of a model’s robustness, generalization, and practical utility.
@@ 2.2 The Need for Efficient VLAs
The primary source of inefficiency lies in the computational footprint of the large VLM backbones themselves. VLA foundational models, characterized by their complex transformer-based architectures and enormous parameter counts (e.g., billions to tens of billions [52, 53, 54, 55]), demand substantial compute power for inference. As shown in Tab.˜I, this results in high latency and insufficient control frequencies, which are often incompatible with the real-time demands of robotic control in dynamic environments. Furthermore, the high dimensionality of visual inputs exacerbates this issue, as processing lengthy sequences of visual tokens through the large model is computationally intensive.
A second critical challenge is the immense cost associated with the large-scale datasets required for VLA pre-training. The collection of real-world robotic demonstration data, as seen in OXE [38] or BridgeData [40], is inherently time-consuming, expensive, and complex. Similarly, generating massive simulation datasets [44] requires significant computational resources and expert knowledge for scenario design. The subsequent pre-training of models on these large-scale datasets incurs exorbitant computational costs, consuming substantial energy and making the process accessible only to well-resourced institutions.
These profound inefficiencies collectively culminate in a critical deployment bottleneck for applications, particularly in resource-constrained settings such as autonomous vehicles and affordable consumer robotics. Specifically, high inference latency, excessive energy consumption, and prohibitive development cost obstruct widespread deployment and the attainment of real-time performance for responsive robotic control. Therefore, the pursuit of efficiency is not merely an optional optimization but a fundamental prerequisite for unlocking the transformative potential of VLAs across the broader robotics landscape. Addressing this imperative, the field of Efficient VLAs has emerged. As illustrated in Fig.˜4, the developmental trajectories of foundational VLAs and efficient VLAs are progressing in parallel. Notably, the timeline reveals a distinct acceleration in the emergence of efficient VLAs, particularly from late 2024 onwards. This proliferation underscores a rapidly intensifying research focus on resource-efficient models, likely driven by the demands of real-world robotic deployment.

### v2

@@ Abstract
Vision-Language-Action models (VLAs) represent a significant frontier in embodied intelligence, aiming to bridge digital knowledge with physical-world interaction. Despite their remarkable performance, foundational VLAs are hindered by the prohibitive computational and data demands inherent to their large-scale architectures. While a surge of recent research has focused on enhancing VLA efficiency, the field lacks a unified framework to consolidate these disparate advancements. To bridge this gap, this survey presents the first comprehensive review of Efficient Vision-Language-Action models (Efficient VLAs) across the entire model-training-data pipeline. Specifically, we introduce a unified taxonomy to systematically organize the disparate efforts in this domain, categorizing current techniques into three core pillars: (1) Efficient Model Design, focusing on efficient architectures and model compression; (2) Efficient Training, which reduces computational burdens during model learning; and (3) Efficient Data Collection, which addresses the bottlenecks in acquiring and utilizing robotic data. Through a critical review of state-of-the-art methods within this framework, this survey not only establishes a foundational reference for the community but also summarizes representative applications, delineates key challenges, and charts a roadmap for future research. We maintain a continuously updated project page to track our latest developments: https://evla-survey.github.io/.
@@ 1 Introduction
Vision-Language-Action Models (VLAs) [67, 10, 187, 77] signify a breakthrough in the evolution of artificial intelligence, endowing interactive physical entities with the capacity for perception and decision-making. These models strategically ground the semantic reasoning of Large Language Models (LLMs) and the cross-modal perception of Vision-Language Models (VLMs) into a unified action space and exhibit unprecedented dexterity and generalizability in complex, unstructured environments. The potential of this paradigm is already manifest across diverse high-stakes domains, ranging from the real-time interaction in autonomous driving [16, 185] and industrial manufacturing [98] to the high-precision requirements of medical robotics [78] and laboratory automation [182].
A synergistic convergence of architectural innovation, sophisticated training regimens, and large-scale data curation propels the rapid maturation of the VLA paradigm. Architectural research focuses on synergizing multimodal backbones [8, 64] with diverse action decoders to facilitate seamless end-to-end mapping from perception to execution, spanning autoregressive policies [67] for sequential prediction to diffusion-based generators [10] for trajectory synthesis. Training methodologies have evolved toward large-scale cross-embodiment pre-training on heterogeneous datasets [9], progressively integrating reinforcement learning [74] and supervised fine-tuning [66] to unlock robust zero-shot generalizability and behavioral dexterity. Meanwhile, data-centric strategies prioritize the systematic acquisition of high-fidelity trajectories through expert demonstrations [12], immersive simulation [19], and generative augmentation [80] to bridge the reality gap and ensure broad generalization across diverse tasks. These multi-dimensional advancements constitute a cohesive research ecosystem that continually pushes the frontiers of autonomous agency in the physical world.
Figure 1: The transition from Foundational VLAs to Efficient VLAs. Foundational VLAs are constrained by inherent challenges, specifically real-time incompatibility, excessive computational cost, and inefficient data collection. Through core methods comprising efficient model design, efficient training, and efficient data collection, efficient VLAs enable high-performance implementation on edge devices.
While these models have demonstrated remarkable generalist capabilities, foundational VLAs inherit the prohibitive computational and data footprints of their scaling-centric LLM and VLM foundations, resulting in a stark mismatch with the stringent latency and power constraints of physical platforms. This discrepancy creates a formidable barrier to transitioning digital intelligence into real-time actuation on edge devices, such as robotic manipulators and mobile agents, manifesting primarily through the following critical bottlenecks: (1) Real-Time Incompatibility: current VLAs suffer from high inference latency and insufficient control frequency [67], making them incompatible with the sub-second control cycles required for responsive and adaptive robotic manipulation; (2) Excessive Computational Cost: the pretraining requirements are prohibitively large, as exemplified by OpenVLA [67] consuming 21,500 A100-GPU hours on a 64-GPU cluster, posing significant barriers to reproducibility and scalability; and (3) Inefficient Data Collection: The reliance on large-scale datasets, with examples like $\pi_{0}$ [10] requiring over 10,000 hours of robotic trajectories, results in a time-consuming and labor-intensive data collection pipeline that severely limits the applicability of such approaches.
To surmount these challenges, the Efficient VLA paradigm has emerged, encompassing resource-conscious embodied models that optimize any efficiency dimension within the model-training-data lifecycle to rectify the suboptimal resource utilization and restricted inference speeds of foundational designs. As illustrated in Fig.˜1, these designs [66, 179, 89, 120, 111, 166] bridge the gap between heavyweight foundations and edge deployment, ensuring high-frequency control and robust performance across diverse robotic scenarios.
@@
Figure 2: The Organization of Our Survey. We systematically categorize efficient VLAs into three core pillars: (1) Efficient Model Design, encompassing efficient architectures and model compression techniques, (2) Efficient Training, covering efficient pre-training and post-training strategies, and (3) Efficient Data Collection, including efficient data collection and augmentation methods. The framework also reviews VLA foundations, key applications, challenges, and future directions, establishing the groundwork for advancing scalable embodied intelligence.
@@
The overarching goal of this survey is to provide a comprehensive view of the technological advances in efficient VLAs. As illustrated in Fig.˜2, we organize the literature in a taxonomy comprising three main categories: (1) Efficient Model Design: encompassing innovative strategies to optimize the architectural and inference efficiency of VLAs. This includes Efficient Architectures, which offers innovative structural optimizations alongside scalable frameworks, and Model Compression, which involves techniques to compress model parameters and refine token handling. (2) Efficient Training: incorporating advanced approaches to reduce the computational and data burdens during VLAs training, containing Efficient Pre-training/Post-training. The former improves efficiency through optimized data and action modeling, while the latter enhances adaptability and performance via customized learning strategies, e.g., supervised fine-tuning and reinforcement learning. (3) Efficient Data Collection: focusing on cutting-edge methods to improve data collection and enrichment for VLAs development, using interactive, simulated, reusable, and self-driven strategies to make dataset acquisition more scalable and efficient. Through this exploration, we aim to provide a comprehensive understanding of the current state of efficient VLAs, thereby highlighting key challenges and advancements in embodied intelligence and robotic learning.
@@
Although several excellent surveys [96, 117, 158, 184, 32, 175] have charted the landscape of general VLAs or the broader field of embodied AI, a systematic review dedicated to the crucial aspect of efficiency in VLAs has been conspicuously absent. This work fills this critical void, uniquely consolidating the disparate research efforts focused on optimizing VLAs for real-world deployment in resource-constrained scenarios, making it an indispensable reference for researchers and practitioners. We hope this survey together with our Github repository can help researchers and practitioners navigate through the literature and serve as a catalyst for inspiring further research on efficient VLAs.
@@
  • Pioneering Survey: To the best of our knowledge, this work presents the first comprehensive survey specifically dedicated to the realm of efficient VLAs that covers the entire “model-training-data” process. It fills a critical gap in the literature and aims to establish a foundational reference for the research community.
@@
The remaining sections of this survey are organized as illustrated in Fig.˜2. We begin in section˜2 by introducing the foundational concepts of Vision-Language-Action models, delineating the computational challenges that necessitate the pursuit of efficiency. Section 3 delves into efficient model design, systematically categorizing techniques that span efficient architectural innovations and model compression. Section 4 investigates efficient training paradigms, encompassing strategies to reduce the computational overhead and dataset demands during both pre-training and post-training stages. Section 5 examines the crucial role of efficient data collection, exploring methodologies for scalable data collection and effective augmentation to maximize data utility. We then survey key applications in section˜6 where efficient VLAs have demonstrated significant potential, highlighting their practical utility in real-world scenarios. Section 7 discusses current challenges and outlines promising future research directions. Finally, section˜8 summarizes the survey.
@@ 2 Vision-Language-Action Models
This section delineates the expansive landscape of VLAs, scrutinizes the critical efficiency bottlenecks in foundational paradigms that necessitate the transition toward efficient VLAs, and underscores the pioneering contribution of this survey as a dedicated roadmap for resource-conscious embodied intelligence.
@@ 2.1.1 Foundational Pipeline
LLM Backbone. After a projector $P(\cdot)$ that ensures seamless alignment and mitigates modality gaps, the multimodal sequence then feeds into the LLM backbone $LLM(\cdot)$—the center for semantic reasoning. Spanning Qwen [3], PaLM [27], Gemma [134], Llama [136], Mamba [45], and Vicuna [26], pre-trained LLMs process the fused embeddings for task planning, which can be formulated as follows:
@@
2.1.2 Datasets
The proliferation of VLAs is fundamentally underpinned by high-fidelity datasets and rigorous benchmarks, which serve respectively as the essential fuel for behavioral emergence and the critical yardstick for performance validation.
@@ 2.1.2 Datasets
Real-world Datasets. A cornerstone of VLA training is the Open X-Embodiment (OXE [104]) dataset. This collaborative effort aggregates data from multiple robotics labs, encompassing a vast array of tasks, robot morphologies, and environments. Its scale and diversity are instrumental for training models with strong generalization and transfer learning capabilities. Complementing OXE are other significant datasets such as BridgeData and BridgeData V2 [38, 139], which focuses on cross-domain tasks, and DROID [65], known for its large-scale dexterous teleoperated data “in the wild”. Expanding this ecosystem, EgoDex [50] introduces high-fidelity egocentric teleoperation data to bridge the gap between human-like perception and dexterous manipulation. Simultaneously, AgiBot-World [12] provides a massive collection of $1M+$ robotic trajectories, specifically curated to enhance the scalability of intelligent agents in complex, real-world industrial and service scenarios. These datasets provide the crucial “embodied” experience necessary for VLAs to learn the dynamics and constraints of the physical world.
Simulation Datasets and Benchmarks. Simulation platforms serve as the strategic engine for synthesizing large-scale, high-fidelity trajectories that circumvent the scalability constraints of physical collection. RoboGen [149] exemplifies the generative simulation paradigm by autonomously creating vast arrays of tasks and demonstrations, while RoboCasa [102] provides a large-scale environment tailored for daily household activities. To address the demand for specialized manipulation, SynGrasp-1B [31] contributes an extensive synthetic collection focusing on diverse grasping primitives, and RoboTwin [19] generates dual-arm cooperative trajectories with diverse data augmentation tricks. These platforms collectively enable the acquisition of diverse experience, facilitating the pre-training of VLAs with enhanced robustness and scalability.
Certain environments bridge the gap between data acquisition and performance validation, serving as both training resources and standardized testbeds. RLBench [54], RoboTwin [19] and Meta-World [172] are widely utilized for their predefined task libraries and consistent evaluation metrics across various manipulation scenarios Beyond data generation, specialized benchmarks have been established to isolate and evaluate specific VLA capabilities. LIBERO [87] focuses on knowledge retention within lifelong learning settings, while CALVIN [99] assesses generalization to novel instructions and long-horizon reasoning. Crucially, SIMPLER [81] provides a high-fidelity evaluation framework designed to bridge the sim-to-real gap, offering a reliable proxy for predicting the real-world performance of VLA policies. Furthermore, VLABench [180] evaluates long-horizon and multi-dimensional reasoning capabilities through massive standard tasks. In summary, these simulated ecosystems constitute a multidimensional evaluation framework that ensures the reliability and practical utility of VLAs before physical deployment.
@@ 2.2 The Need for Efficient VLAs
From an architectural perspective, the primary impediment resides in the massive computational footprint of transformer-based VLM backbones. Characterized by billions of parameters [8, 64, 36, 20], these models incur prohibitive inference latency, resulting in control frequencies that fail to meet the stringent real-time requirements of dynamic robotic actuation, as evidenced in Tab.˜I. This inefficiency is further exacerbated by the high dimensionality of visual inputs, which necessitates the intensive processing of dense token sequences through deep layers.
Regarding training methodologies, the current paradigm is burdened by the exorbitant computational budgets and energy consumption required for large-scale pre-training. Achieving robust generalization across heterogeneous embodiments demands massive compute resources, creating a significant barrier to entry and restricting the iterative development of VLAs to well-resourced institutions.
Lastly, from a data perspective, the acquisition of high-fidelity trajectories remains a critical bottleneck. Whether through labor-intensive real-world demonstrations [104, 139] or expert-intensive simulation design [149], the curation of diverse datasets is inherently time-consuming and difficult to scale, significantly inflating the overall cost of the development lifecycle.
These profound inefficiencies collectively culminate in a critical deployment bottleneck for applications, particularly in resource-constrained settings such as autonomous vehicles and affordable consumer robotics. Therefore, the pursuit of efficiency is not merely an optional optimization but a fundamental prerequisite for unlocking the transformative potential of VLAs across the broader robotics landscape. Addressing this imperative, the field of Efficient VLAs has emerged. As illustrated in Fig.˜4, the developmental trajectories of foundational VLAs and efficient VLAs are progressing in parallel. Notably, the timeline reveals a distinct acceleration in the emergence of efficient VLAs, particularly from late 2024 onwards. This proliferation underscores a rapidly intensifying research focus on resource-efficient models, likely driven by the demands of real-world robotic deployment.